# Large-charge Rényi entropy

### v1

@@ IV.2 Large-charge fluctuation probability
$\log Z(\beta,Q)=-\beta E_{\rm g.s.}(Q)-\sum_{m,k}\log\left(1-e^{-\beta\omega_{m,k}}\right).$ (14)
@@
$\log p(Q)=-2\pi{V_{2}}\left(c_{3/2}\,q^{\frac{3}{2}}+c_{1/2}\,q^{\frac{1}{2}}+c_{0}+O(q^{-\frac{1}{2}})\right).$ (15)
@@
Note that the result is intuitive in that the probability of measuring charge $Q$ inside a ball-shaped region in vacuum is exponentially damped at large-$Q$. Furthermore, the relation $p(Q)=O(e^{-q^{3/2}})$ is presumably universal for any three-dimensional theories with a proper thermodynamic limit because of dimensional analysis [2]. This is true even when they are not described by (9) at large-charge, e.g., the free Fermion theory [20].
IV.3 Large-charge entanglement entropy
The charged Rényi entropy can be written using the charged thermal partition function, as shown in (4),
$S_{n}(Q)\equiv\frac{\log Z(2\pi Rn,Q)-n\log Z(2\pi R,Q)}{1-n}.$ (16)
Thus it is immediate that the vacuum contribution in (14) cancels out, and that it depends only on the one-particle spectrum, $\omega_{m,k}$. It is also apparent by now that $S_{n}(Q)$ has an expansion in terms of $1/q$, and that its leading order is $O(1)$, which simply comes from the fluctuation Lagrangian (10). Interestingly, this number is universal across different CFTs as long as they are in the large-charge universality class described by (9); The only theory dependence lies in subleading corrections.
We now numerically compute the leading $O(1)$ part of $S_{n}(Q)$, denoted as $S_{n}^{(0)}(Q)$, using the numerically obtained values of $\omega_{m,k}$. As an example, we show a plot of $S_{1}^{(0)}(Q)$ as a function of $u_{0}$ in Figure 1; we see that it behaves as $O(e^{u_{0}})$ at large $u_{0}$, which continues to be true for other values of $n$. By fitting the numerical value of $S_{n}^{(0)}(Q)$ against an expansion in terms of $V_{2}=2\pi\left(\frac{e^{u_{0}}}{2}-1+\frac{e^{-u_{0}}}{2}\right)$, we see that
$S_{n}(Q)=\left(A_{n}\pi^{-1}V_{2}+O({V_{2}}^{0})\right)q^{0}+O(q^{-1/2}),$ (17)
@@ IV.3 Large-charge entanglement entropy
$A_{1}=0.071,\quad A_{2}=0.031,\quad A_{3}=0.024.$ (18)
Incidentally, $A_{n}$ is approximated well by $A_{n}\approx 0.018+0.052n^{-2}$, but its theoretical justification is unclear. To summarise, the charged Rényi entropy in the large-charge universality described by (9) converges to a universal $O(1)$ number at large-charge, as given in (17).
Figure 1: A plot of the large-charge von Neumann entropy $S_{1}^{(0)}(Q)$ in terms of $u_{0}$ (red dots) with a comparison to a fit (blue line). We see a perfect fit at large $e^{u_{0}}$.
@@ V Conclusions and Outlook
We can also study various other limits of the symmetry-resolved entanglement entropy. For example, we can take the large-representation limit in a theory with a non-Abelian symmetry. This would result in various one-particle excitations containing non-relativistic NG bosons as well as phonons [3, 42, 43, 44, 45]. Another direction would be to study the limit of large spin, which is known to be an interesting solvable limit for operator dimensions and the OPE coefficients [46, 47, 48, 49, 50]. It would be interesting to ask if it also translates to the simplification of the modular Hamiltonian spectrum. One can also study the regime of large-charge and large-spin, which is known to exhibit various interesting phases with e.g., vortex lattices and the giant vortex [51, 52, 53, 54, 41].
@@ Acknowledgements
The author thanks Simeon Hellerman, Kantaro Ohmori, and Adar Sharon for valuable discussions. This work is supported by a Grant-in-Aid for JSPS Fellows No. 22KJ1777, a Grant-in-Aid for Early-Career Scientists No. 25K17387, and by a MEXT KAKENHI Grant No. 24H00957.

### v2

@@ 4.3 Large-charge fluctuation probability
$\log Z(\beta,Q)=-\beta E_{\rm g.s.}(Q)-\sum_{m,k}\log\left(1-e^{-\beta\omega_{m,k}}\right).$ (4.5)
@@
$\log p(Q)=-2\pi{V_{2}}\left(c_{3/2}\,q^{\frac{3}{2}}+c_{1/2}\,q^{\frac{1}{2}}+c_{0}+O(q^{-\frac{1}{2}})\right).$ (4.6)
@@
Note that the result is intuitive in that the probability of measuring charge $Q$ inside a ball-shaped region in vacuum is exponentially damped at large-$Q$. Furthermore, the relation $p(Q)=O(e^{-q^{3/2}})$ is presumably universal for any three-dimensional theories with a proper thermodynamic limit because of dimensional analysis [2]. This is true even when they are not described by (3.1) at large-charge, e.g., the free Fermion theory [24].
4.4 Large-charge entanglement entropy
The charged Rényi entropy can be written using the charged thermal partition function, as shown in (2.4),
$S_{n}(Q)\equiv\frac{\log Z(2\pi Rn,Q)-n\log Z(2\pi R,Q)}{1-n}.$ (4.7)
Thus it is immediate that the vacuum contribution in (4.5) cancels out, and that it depends only on the one-particle spectrum, $\omega_{m,k}$. It is also apparent by now that $S_{n}(Q)$ has an expansion in terms of $1/q$, and that its leading order is $O(1)$, which simply comes from the fluctuation Lagrangian (4.1). Interestingly, this number is universal across different CFTs as long as they are in the large-charge universality class described by (3.1); The only theory dependence lies in subleading corrections.
We now numerically compute the leading $O(1)$ part of $S_{n}(Q)$, denoted as $S_{n}^{(0)}(Q)$, using the numerically obtained values of $\omega_{m,k}$. As an example, we show a plot of $S_{n=1}^{(0)}(Q)$ as a function of $u_{0}$ in Figure 1; we see that it behaves as $O(e^{u_{0}})$ at large $u_{0}$, which continues to be true for other values of nn. [^8] This is by no means trivial. With a physical boundary condition imposed on the hyperbolic space, there is no guarantee that the Rényi entropy scales with the regularised volume of the hyperbolic space, as is usually claimed.
By fitting the numerical value of $S_{n}^{(0)}(Q)$ against an expansion in terms of $V_{2}=2\pi\left(\frac{e^{u_{0}}}{2}-1+\frac{e^{-u_{0}}}{2}\right)=2\pi\frac{R}{\delta}+O(1)$, we see that [^9] Note that this is not the usual expansion where $V_{2}$ is the largest parameter. in which the $O(1)$ part is commonly argued to be universal.
$S_{n}(Q)=\left(A_{n}\pi^{-1}V_{2}+O({V_{2}}^{0})\right)q^{0}+O(q^{-1/2}),$ (4.8)
@@ 4.4 Large-charge entanglement entropy
$A_{1}=0.071,\qquad A_{2}=0.031,\qquad A_{3}=0.024.$ (4.9)
To reiterate, these numbers are universal for any CFTs in our large-charge universality class.
Incidentally, $A_{n}$ is approximated well by $A_{n}\approx 0.052n^{-2}+0.018$ at small $0.5\lesssim n\lesssim 5$. To see why, notice that (4.1) is that of a minimally coupled free scalar, and so its high-temperature thermal partition function can be approximated by that of a conformal scalar with corrections that go as $O(\beta)$. This allows us to use a general analysis of high-temperature limit of CFTs using thermal EFTs [25, 26], which leads to $S_{n}=f/n^{2}+g/n+\text{(const.)}$, consistent with our fit. Note that $g$ is related to the boundary entropy and so its absence could be explained by the use of the Neumann boundary condition. [^10] We thank Sridip Pal for discussions on this paragraph.
Figure 1: A plot of the large-charge von Neumann entropy $S_{n=1}^{(0)}(Q)$ in terms of $u_{0}$ (red dots) with a comparison to a fit (blue line). We see a perfect fit at large $e^{u_{0}}$.
@@ 5 Conclusions and Outlook
We can also study various other limits of the symmetry-resolved entanglement entropy, or other interesting symmetry-enriched entanglement quantities such as entanglement asymmetry [48, 49, 50]. For example, we can take the large-representation limit in a theory with a non-Abelian symmetry. This would result in various one-particle excitations containing non-relativistic NG bosons as well as phonons [3, 51, 52, 53, 54]. Another direction would be to study the limit of large spin, which is known to be an interesting solvable limit for operator dimensions and the OPE coefficients [55, 56, 57, 58, 59, 60, 61]. It would be interesting to ask if it also translates to the simplification of the modular Hamiltonian spectrum. One can also study the regime of large-charge and large-spin, which is known to exhibit various interesting phases with e.g., vortex lattices and the giant vortex [62, 63, 64, 65, 47].
@@ Acknowledgements
The author thanks Simeon Hellerman, Kantaro Ohmori, Sridip Pal, and Adar Sharon for valuable discussions. This work is supported by a Grant-in-Aid for JSPS Fellows No. 22KJ1777, a Grant-in-Aid for Early-Career Scientists No. 25K17387, and by a MEXT KAKENHI Grant No. 24H00957.